### v2

# Safety-Aware Multi-Agent Learning for Dynamic Network Bridging

Abstract
Addressing complex cooperative tasks in safety-critical environments poses significant challenges for multi-agent systems, especially under conditions of partial observability. We focus on a dynamic network bridging task, where agents must learn to maintain a communication path between two moving targets. To ensure safety during training and deployment, we integrate a control-theoretic safety filter that enforces collision avoidance through local setpoint updates. We develop and evaluate multi-agent reinforcement learning safety-informed message passing, showing that encoding safety filter activations as edge-level features improves coordination. The results suggest that local safety enforcement and decentralized learning can be effectively combined in distributed multi-agent tasks.
I Introduction
Despite the promising performance of multi-agent reinforcement learning (MARL) in simulations, its application in real-world, safety-critical scenarios raises significant concerns. A key challenge is collision avoidance, especially in applications like autonomous driving, where failure to prevent collisions could result in severe consequences. Moreover, the inherent nature of multi-agent systems (MAS) often involves partial observability, further increasing the complexity of these problems by limiting the information available to each agent about the environment and the states of other agents.
This challenge highlights a fundamental limitation of all reinforcement learning (RL)-based approaches: while agents learn to optimize predefined reward functions, their behavior lacks guarantees and remains inherently unpredictable. The sole reliance on MARL’s reward functions to ensure safety is insufficient [1]. Even with extensive training and monitoring of an agent’s performance based on expected rewards, or other performance measures, it is impossible to exhaustively examine all potential scenarios in which the agent might fail to act safely or predictably.
In ensuring safety among agents, various control-theoretic approaches and hybrid learning approaches have been developed, ranging from Model Predictive Control (MPC) [2] to Reference Governor (RG) techniques [3], and from Control Barrier Functions (CBFs) [4] to Reachability Analysis (RA) methods [5]. In robotics applications, those methods are known as safety filters [6], and they are deployed to ensure safety regardless of the algorithm used for the agent to accomplish a task. While safety filters can effectively prevent unsafe behavior, they are often conservative by design. This conservatism may lead to conflicts with task objectives, especially in MAS where coordination is essential.
In this work, we develop a training framework for MARL that incorporates safety constraints during policy learning. We adapt an optimization-free safety tracking control algorithm developed in [7], which is based on ellipsoidal positively invariant sets. We show how this safety filter can be adapted to the context of MAS and how safety is guaranteed. We integrate this type of filter in the agents’ training process, allowing them to learn decentralized policies and avoid situations where the filter compromises task objectives.
Figure 1: Decentralized swarm coordination for dynamic network bridging
We demonstrate our approach in the dynamic network bridging task [8] (Fig. 1) where a number of agents establish and maintain a communication path between two moving targets. Our key contributions are as follows: 1) A novel decentralized control framework that integrates a safety filter based on invariant sets into the MARL agents’ training process, and restricts the effect of each agent’s movement updates to only its one-hop neighbors, enabling efficient local coordination; 2) We augment the learning process with safety-informed message passing, where agents observe safety filter activations as edge-level features during training, leading to improved coordination 3) An algorithm that updates setpoints while preserving safety conditions through communication with affected neighbors; 3) An analytical computationally tractable condition verifying potential safety violations during updates. Overall our approach provides effective coordination among agents while ensuring safety constraints.
II Preliminaries
II-A Task description
Dynamic network bridging is a multi-agent task focused on addressing the problem of dynamically establishing and maintaining a communication link between two mobile, independent nodes (targets) using a swarm of autonomous agents taking decentralized decisions [8]. Agents and targets are moving in a 2D grid space and have a limited communication range, within which they exchange information. This region can be defined as a 2D ball of radius $r$, centered at its position $p_{i}(t)=[p_{x,i}(t),p_{y,i}(t),p_{z,i}(t)]^{T}$. Based on this definition, two agents $i$ and $j$ can communicate if and only if their regions intersect:
$\mathcal{B}_{r}(p_{i}(t))\cap\mathcal{B}_{r}(p_{j}(t))\neq\emptyset.$ (1)
where
$\mathcal{B}_{r}(p_{i}(t))=\left\{p\in\mathbb{R}^{3}:\|p-p_{i}(t)\|^{2}\leq r^{2}\right\}$
In our setup, we consider $p_{z,i}$ fixed, therefore the one-hop communication region can be described by a 2D ball. Formally, the system is represented as a dynamic graph $\mathcal{G}(t)=(\mathcal{V},\mathcal{E}(t))$ with nodes $\mathcal{V}$ corresponding to agents and targets. An edge $(u,v)$ exists between nodes if their Euclidean distance at time-step $t$ is within a predefined communication range.
The task reflects scenarios where communication infrastructure is absent or unreliable, such as in disaster zones or remote environments. Agents must position themselves to form an ad-hoc, decentralized communication network that connects the two mobile targets. Each agent operates based on local observations and limited-range communication, with no access to global state information.
II-B Multi-agent reinforcement learning
In a typical multi-agent cooperative scenario, $N$ agents interact within an environment to complete a task. Noisy and limited sensors may prevent each agent from observing the full state of the environment, allowing access only to partial observations. With partial observability, an agent no longer knows the true state of the environment, but instead needs to maintain a belief state - a probability distribution over states estimated from its observation history - which it uses to select actions. Decision-making processes in such contexts can be formulated as a decentralized partially observable Markov decision process (Dec-POMDP) [9, 10].
A Dec-POMDP is defined by a tuple $(S,\mathbf{A},\mathcal{O},T,R,\Omega,\gamma,N)$ where $S$ is a finite set of states of the environment, $\mathbf{A}$ is a set of joint action spaces, and $N$ is the number of agents, $T$ is the state transition probability function which gives the probability of transitioning from state $s$ to state $s^{\prime}$ when joint action $\mathbf{a}=(a_{1},\cdots,a_{N})$ is taken; $R$ is the joint reward function that maps states and joint actions to real numbers and is used to specify the goal of the agents; $\Omega$ is a set of joint observations $O_{i}$, where $O_{i}$ is the observation set for agent $i$; $\mathcal{O}$ is the joint observation probability function, which gives the probability of receiving joint observation $\mathbf{o}=(o_{1},\cdots,o_{N})$ after taking joint action $\mathbf{a}$ and ending up in state $s^{\prime}$; and $\gamma$ is the reward discount factor, $0\leq\gamma\leq 1$. The goal of the agents is to maximize a joint reward function. Solving a Dec-POMDP optimally is computationally intractable, therefore, various approximate solution methods [11, 12] have been used, such as heuristic search, dynamic programming, value or policy-based approximators with neural networks etc.
II-C Distributed autonomous swarm formation for dynamic network bridging
We formulate our problem as a Dec-POMDP, where each agent has restricted sensing capabilities given by a specific communication range, leading to incomplete information about the global state of the system outside such area. However, whenever two entities (agents or targets) are within range, a communication channel becomes available for them to exchange information. This channel enables the possibility for agents to know each other features and execute distributed strategies requiring message exchange to coordinate. At the beginning of every time-step, agents make decentralized decisions, generating target points $w$, utilizing their local observation to choose the direction of their next movement. To solve the Dec-POMDP that corresponds to our task we used a graph neural network (GNN) approximator that encodes the graph structure of the problem.
II-D Agent Dynamics
We consider a linear time-invariant (LTI) of the form
$\dot{x}=A(x-x_{sp})$ (2)
where $x\in\mathbb{R}^{n}$ is the state of the system, and the origin is translated to the setpoint $x_{sp}\in\mathbb{R}^{n}$. The matrix $A\in\mathbb{R}^{n\times n}$ is Hurwitz (i.e., $\mathrm{Re}(\lambda_{i})<0$), where $\lambda_{i}\in\mathbb{C}$ is the $i$-th eigenvalue of $A$. The matrix $A$ can be seen as the closed-loop system matrix obtained by applying a state-feedback controller (i.e., $u=-K(x-x_{sp})$), where $u$ is the control input.
For the asymptotically stable LTI system (2), there exists a quadratic Lyapunov function of the form
$V(x)=(x-x_{sp})^{T}P(x-x_{sp})$ (3)
where $P$ is a symmetric positive definite matrix that satisfies the Lyapunov equation
$A^{T}P+PA=-Q$ (4)
for a given symmetric positive definite matrix $Q>0$.
Defining the $P$-norm as
$\lVert x-x_{sp}\rVert_{P}\triangleq\sqrt{(x-x_{sp})^{T}P(x-x_{sp})},$ (5)
for a fixed value $c>0$, the Lyapunov function (3) defines an ellipsoid, which can be written in terms of (5) as follows:
$\mathcal{E}_{c}(x_{sp})=\{x\in\mathbb{R}^{n}\mid\lVert x-x_{sp}\rVert_{P}^{2}\leq c\}.$ (6)
From (4), $V(x(t))$ is a monotonically decreasing function; therefore, for any $x(t)$ starting within the ellipsoid (6), the trajectory never leaves it. Hence, $\mathcal{E}_{c}$ is called a positively invariant set [13]. For the rest of the paper, we will refer to the state of the agent $i$ as $x_{i}$ and for the related setpoint as $x_{sp,i}$. In this context, the tracking algorithm updates the setpoint of the agent to reach a desired position. This formulation can be extended to more general nonlinear systems by using linearization [13], or Lyapunov-like methods such as quadratic boundedness [14], as used in the context of UAVs in [7], where ellipsoidal positively invariant sets are provided. We use this last case in our experiments.
II-E Safety
Definition 1
Given two agents $i$ and $j$, with distinct equilibrium points $x_{sp,i}\neq x_{sp,j}$ and initial states $x_{i}(t_{0})\in\mathcal{E}_{c}(x_{sp,i})$ and $x_{j}(t_{0})\in\mathcal{E}_{c}(x_{sp,j})$, respectively, agents $i$ and $j$ are deemed safe relative to each other if:
$\mathcal{E}_{c}(x_{sp,i})\cap\mathcal{E}_{c}(x_{sp,j})=\emptyset.$ (7)
This definition leverages the invariant properties of (6), ensuring agents’ trajectories remain confined within their respective safety ellipsoids, thereby preventing collisions.
To operationalize this safety criterion within the two-dimensional communication space, we impose the following:
Assumption 1
The projection of $\mathcal{E}_{c}(x_{sp,i})$ onto the 2-D plane defined by the coordinates $(p_{x},p_{y})$ must always lie within the 2-D communication ball $\mathcal{B}_{r}(p_{i}(t))$, where $p_{i}(t)=[p_{x,i}(t),p_{y,i}(t)]$ represents the actual position of agent $i$.
III Safety Filter for Tracking Control
The proposed algorithm consists of two steps: first, computing a feasible safe step (Figure 2(a)); and second, checking whether the new position would interfere with another agent (Figure 2(b)).
Let us start with the first step which has been borrowed from [7]. For each agent, we select two scalars $c$ and $s$ such that $0<s<c$ therefore $\mathcal{E}_{s}(x_{sp,i})\subset\mathcal{E}_{c}(x_{sp,i})$.
Definition 2
We can say that the agent has reached the setpoint $x_{sp,i}$ if and only if $x_{i}(t)\in\mathcal{E}_{s}(x_{sp,i})$.
Assuming that the agent $i$ is moving from the target point $w_{i}$ to the target point $w_{i+1}$, generated by its decentralized policy, the idea is to generate a sequence of setpoints on the line that connects the two target points. Once the state of agent $i$, $x_{i}$ reached the setpoint $x_{sp,i}$, a new setpoint $x^{\prime}_{sp,i}$ is generated following the rule
$x^{\prime}_{sp,i}=x_{sp,i}+(\sqrt{c}-\sqrt{s})v$ (8)
where $v\in\mathbb{R}^{n}$ with $\|v\|_{2}=1$, indicating the direction of the line connecting the two target points. As showed in [7]
$x_{i}(t)\in\mathcal{E}_{s}(x_{sp,i})\Rightarrow x_{i}(t)\in\mathcal{E}_{c}(x^{\prime}_{sp,i}).$ (9)
Fig. 2(a) illustrates the process of updating the setpoint, presented in 2-D space for clearer visualization, though it actually corresponds to a state space with n=12 components, as applicable to quadrotors.
(a) Setpoint update for agent $i$ in the state space.
(b) $m_{\lambda}$ is on the segment that connects the two setpoints.
Figure 2: Safe tracking control and safety verification.
Thanks to (8) and Assumption 1, the projection of $\mathcal{E}_{c}(x^{\prime}_{sp,i})$ is contained in $\mathcal{B}(p_{i}(t))$.
Proposition 1
Considering agent $i$ moving from target point $w_{i}$ to $w_{i+1}$, and updating the setpoint $x_{sp,i}$ every time $x_{i}(t)\in\mathcal{E}_{s}(x_{sp,i})$ by using (8). Considering also Assumption 1 true, then the safety condition (7) can be violated only with all the agents $j\neq i$ that satisfy (1) (i.e. within the one-hop communication range).
Proof:
From Assumption 1, the ball $\mathcal{B}_{r}(p_{i}(t))$ encompasses the projection onto the 2-D space of $\mathcal{E}_{c}(x_{sp,i})$ for any current state $x_{i}(t)\in\mathcal{E}_{c}(x_{sp,i})$. Utilizing (8) to adjust the setpoint ensures that $x_{i}(t)$ remains within $\mathcal{E}_{c}(x^{\prime}_{sp,i})$, signifying that its 2-D space projection still falls within $\mathcal{B}_{r}(p_{i}(t))$. Thus, should the updated setpoint $x^{\prime}_{sp,i}$ contravene (7), such a violation would only occur concerning agents situated within the one-hop communication range. ∎
Proposition 1 highlights a critical aspect of setpoint updates: setpoint updates only impact safety for agents within communication range. This ensures agent $i$ can exchange messages with affected neighbors before executing updates. This allows coordinating to preemptively address any potential safety concerns from the new setpoint through direct communication with impacted agents.
Building upon Proposition 1, let’s define $\mathcal{D}_{i}(t)$ as the set comprising all agents $j\neq i$ that are within the one-hop communication range of agent $i$ at time $t$. The safe setpoint update algorithm for each agent $i$ is given in Algorithm 1. Once agent $i$’s state $x_{i}(t)$ reaches its current setpoint $x_{sp,i}$ (as per Definition 2), it computes a new target setpoint $x^{\prime}_{sp,i}$. Before updating, agent $i$ checks for any possible safety violations by evaluating the safety condition (7) for all agents in $\mathcal{D}_{i}(t)$. If no violation is detected, $x_{sp,i}$ is updated to the new $x^{\prime}_{sp,i}$; otherwise, the previous setpoint is retained.
Input: Agent $i$ with setpoint $x_{sp,i}$, $\mathcal{D}_{i}(t)$
Output: Updated setpoint $x_{sp,i}$
potential_safety_violation $\leftarrow$ 0;
if $x_{i}(t)\in\mathcal{E}_{s}(x_{sp,i})$ then
       Compute $x^{\prime}_{sp}\leftarrow$ (8);
       for $j\in\mathcal{D}_{i}(t)$ do
             if $\mathcal{E}_{c}(x_{sp,i})\cap\mathcal{E}_{c}(x_{sp,j})\neq\emptyset$  then
                   potential_safety_violation $\leftarrow 1$;
                   break;
             end if
       end for
       if potential_safety_violation ==absent$=$ = 00 then
             $x_{sp,i}\leftarrow x^{\prime}_{sp,i}$;
       end if
end if
Algorithm 1 Setpoint Update Algorithm
By applying Algorithm 1 to each agent, we ensure that the safety condition (7) is satisfied throughout the system. This is guaranteed by the fact that an agent can only update its setpoint after verifying that the new target does not violate safety for any agents within its one-hop communication range, as stated in Proposition 1. The proof of safety relies on the following assumptions:
Assumption 2
The initial setpoints of all agents satisfy the safety condition (7).
Assumption 3
During the setpoint update for agent $i$, all agents $j\in\mathcal{D}_{i}(t)$ cannot change their setpoints $x_{sp,j}$.
Proposition 2
Considering the MAS described where the agents are LTI systems (2) and Assumptions 1, 2 and 3, then applying Algorithm 1 to each agent guarantees that the safety condition (7) is satisfied for all agents.
Proof:
By Assumption 2, the initial setpoints of all agents satisfy the safety condition (7). Whenever an agent $i$ updates its setpoint using Algorithm 1, it checks if the new target setpoint $x_{sp,i}^{\prime}$ violates the safety condition for any agent $j\in\mathcal{D}_{i}(t)$. If a violation is detected, the setpoint is not updated, ensuring that the current safe solution is maintained. Proposition 1 guarantees that safety is only impacted for agents within the one-hop communication range, which agent $i$ can directly communicate with before executing the update. Furthermore, Assumption 3 prevents agents in $\mathcal{D}_{i}(t)$ from changing their setpoints during agent $i$ ’s update, ensuring a consistent safety evaluation. Therefore, by applying Algorithm 1 to each agent, the safety condition (7) is upheld throughout the system’s operation. ∎
In the second step, we need to verify whether the safety condition (7) is going to be violated. Our objective is to demonstrate that if (7) is violated, then there exists a point on the line connecting $x_{sp,i}$ and $x_{sp,j}$, inside the overlap region of two ellipsoids $\mathcal{E}_{c}(x_{sp,i})$ and $\mathcal{E}_{c}(x_{sp,j})$ (Fig. 2(b)). This overlap region, denoted by $\mathcal{E}_{c}(x{sp,i})\cap\mathcal{E}_{c}(x_{sp,j})$, satisfies the following condition:
$\lambda(x-x_{sp,i})^{T}P(x-x_{sp,i})+(\lambda-1)(x-x_{sp,j})^{T}P(x-x_{sp,j})\leq c$ (10)
for $0\leq\lambda\leq 1$ with $x_{sp,i}\neq x_{sp,j}$.
Proposition 3
Let us consider the ellipsoids $\mathcal{E}_{c}(x_{sp,i})$ and $\mathcal{E}_{c}(x_{sp,j})$ defined as in (6). The set of points $x$ satisfying (10) for all $\lambda\in[0,1]$ is either empty, a single point, or an ellipsoid:
^$\hat{\mathcal{E}}_{K_{\lambda}}(m_{\lambda})=\left\{x\in\mathbb{R}^{n}:(x-m_{\lambda})^{T}P^{-1}(x-m_{\lambda})\leq K_{\lambda}\right\}$ (11)
where
$m_{\lambda}$ $=\lambda x_{sp,i}+(1-\lambda)x_{sp,j}$
$K_{\lambda}$ $=1-\lambda(1-\lambda)(x_{sp,j}-x_{sp,i})^{T}P(x_{sp,j}-x_{sp,i}).$
Proof:
See [15]. ∎
Proposition 3 shows that if there is an ellipsoid violation, then the point $m_{\lambda}$ belongs to the intersection Fig. 2(b):
$\mathcal{E}_{c}(x_{sp,i})\cap\mathcal{E}_{c}(x_{sp,j})\neq\emptyset\rightarrow m_{\lambda}\in\mathcal{E}_{c}(x_{sp,i})\cap\mathcal{E}_{c}(x_{sp,j})$
Therefore, to test if the two ellipsoids intersect we need to find if there exists at least one point on the segment joining $x_{sp,i}$ and $x_{sp,j}$ that satisfies (10). To do so we consider both ellipsoids
$\left\{\ \begin{array}[]{l}(x-x_{sp,i})^{T}P(x-x_{sp,i})\leq\rho\\ (x-x_{sp,j})^{T}P(x-x_{sp,j})\leq\rho\\ \end{array}\right.$ (12)
We replace $x$ with $m_{\lambda}$ since we are checking only a point in the segment joining the two setpoints. Defining $d\triangleq x_{sp,j}-x_{sp,i}$, (12) can be rewritten as
$\left\{\ \begin{array}[]{r}\lambda^{2}d^{T}Pd\leq\rho\\ (\lambda-1)^{2}d^{T}Pd\leq\rho\\ \end{array}\right.\Rightarrow\left\{\ \begin{array}[]{r}\lambda^{2}\|d\|_{P}^{2}\leq\rho\\ (\lambda-1)^{2}\|d\|_{P}^{2}\leq\rho\\ \end{array}\right.$ (13)
Solving for $\lambda$ we obtain
$\left\{\ \begin{array}[]{l}\lambda^{2}\leq\frac{c}{\|d\|_{P}^{2}}\\ \lambda^{2}\|d\|_{P}^{2}-2\lambda\|d\|_{P}^{2}+\|d\|_{P}^{2}\leq c\\ \end{array}\right.\Rightarrow 1-2\sqrt{\frac{c}{\|d\|_{P}^{2}}}\leq 0$ (14)
The two ellipsoids $\mathcal{E}_{c}(x_{sp,i})$ and $\mathcal{E}_{c}(x_{sp,j})$ intersect if
$2\sqrt{\frac{c}{\|d\|_{P}^{2}}}\geq 1\rightarrow\|d\|_{P}^{2}\leq 4c$ (15)
Remark 1
Although Algorithm 1 ensures safety according to Proposition 2, it does not guarantee target achievement. Agents within one-hop communication range may become immobilized at a certain position because any attempt to reach the target points could potentially violate the safety condition (7). We intentionally refrain from employing any enforcer to resolve deadlock scenarios. Instead, we allow the agents to learn how to avoid or recover from them through the training process.
In summary, our analysis provides the following key theoretical results:
  • Each agent’s setpoint update only impacts the safety of neighboring agents within its one-hop communication range. This localized effect is an important property that enables decentralized coordination.
  • we establish that Algorithm 1 guarantees the preservation of the safety condition in (7) for all agents,
  • We provide (15) a computationally tractable method for evaluating possible safety condition (7) violations.
Together, these theoretical results form the basis for our safe and distributed multi-agent control framework, ensuring that agents can dynamically update their setpoints while maintaining the prescribed safety guarantees through local communication and coordination.
Training Conditions Evaluation Results
Safety Filter Explicit Penalty Truncation Avg Coverage Avg Safety Interventions
Baseline No No No $(37.63\pm 16.62)\%$ $26.99\pm 12.31$
Approach A Yes No No $(50.14\pm 12.72)\%$ $20.45\pm 7.49$
Approach B Yes Yes No $(45.02\pm 11.81)\%$ $15.25\pm 7.82$
Approach C Yes Yes Yes $(10.86\pm 6.44)\%$ $9.76\pm 5.72$
TABLE I: Comparison of the trained agents performance across 100 evaluation episodes.
IV Decentralized Safety-Aware Agents
This section details the components of our approach to ensure safe interactions while learning effective positioning strategies in networked and decentralized multi-agent tasks, by combining MARL with our safety filter mechanism. The core intuition is to employ Algorithm 1 since the training phase and inform the agents of eventual safety filter activations.
IV-A Guiding learning via decentralized safety filters
By enforcing adherence to explicit safety conditions defined in (7), agents experience realistic operational constraints and encounter possible deadlock scenarios whenever direct trajectories toward their target points trigger the safety filter. These constraints inherently shape the policy optimization landscape during training: 1) creating deadlock conditions agents might be blocked by their respective safety filters unless they reconsider their actions; 2) avoiding unsafe states of the environment, which are not going to be visited by the agents. Consequently, the agents need to leverage interactions and communication with neighboring agents to collaboratively avoid or resolve deadlock situations by generating alternative, safety-compliant paths to achieve their objectives effectively.
IV-B Safety-informed edges
Leveraging the graph nature of the task introduced in Section II-C, we explicitly introduce edge-level features linked to the conditions determining the activation of the decentralized safety filter. Specifically, edge features comprise base attributes controlling the activation of the safety filter and a binary safety activation indicator between connected agents. Formally, given two connected agents $i$ and $j$, the edge features are defined as:
$f_{l}^{ij}=[\mathbf{f},S_{i,j}],$ (16)
where $\mathbf{f}$ denotes a vector of available attributes related to the activation of the decentralized safety filter, and the binary indicator $S_{i,j}$ explicitly signals whether the safety filter has been activated between the two agents, thereby preventing setpoint update.
IV-C Policy parameterization
To demonstrate our approach we train the agents with a simple independent Q-Learning algorithm [16] and employ the same GNN architecture proposed in [8] for action-value estimation. The architecture integrates Graph Attention Network (GAT) [17] and Long Short-Term Memory (LSTM) layers [18], enabling effective modeling of spatial and temporal dependencies essential for collaborative decision-making. The GAT layers allow agents to dynamically attend to and integrate neighbor information, significantly improving inter-agent coordination by adaptively sharing latent representations through message passing. The LSTM captures temporal dependencies and partial observability, allowing agents to leverage historical context in their decisions. The architecture is trained using a centralized training decentralized execution (CTDE) framework, where the agents optimize the same action-value function parameterization during training while acting independently based on local observations and the interaction with the other agents.
V Experimental Setup
In this section, we describe in detail the specific environment implementation and training configurations employed for our MARL agents.
V-A Dynamic network bridging environment implementation
Environment: The simulated environment is normalized to a $[0,1]\times[0,1]$ grid, though agents can move beyond these limits. Each entity has a communication radius of 0.20, which defines the area within which it can detect and communicate with others.
Agent observations: At each time step, agents receive an observation consisting of their local one-hop graph structure. This includes node-level features for themselves and all directly connected neighbors, as well as edge-level features for each neighbor pair. An example is shown in Figure 3, where Agent A observes Agent B. The observation includes both their node features $f_{n}^{A},f_{n}^{B}$ and the corresponding edge features $f_{i}^{AB}$ from a successful connection. Node-level features include the agent’s ID, current coordinates, last action taken, and the positions of the targets A and B. These features provide basic local context for decision-making. Edge-level features reflect spatial and safety-related information, as introduced in Section IV-B. For each neighboring agent $j$, agent $i$ observes relative position offsets $\Delta p_{x}=p_{x,i}-p_{x,j}$, $\Delta p_{y}=p_{y,i}-p_{y,j}$, and their Euclidean distance $d_{i,j}=\sqrt{\Delta p_{x}^{2}+\Delta p_{y}^{2}}$. Including safety-informed edge features allows agents to adjust their actions based not only on position but also on the state of the safety filter. This supports more informed and coordinated decision-making during training and execution.
Agent actions: At the beginning of every time-step, agents make decentralized decisions, generating target points $w$, utilizing their local observation to choose the direction of their next movement. This is encoded in a two-dimensional action space, with each dimension having three options: move forward, move backward, or hold along the corresponding $p_{x}$ or $p_{y}$ axis. Given the direction chosen by the agents, a new target point $x_{sp}$ is calculated for each agent using a fixed offset equal for both the $p_{x}$ and $p_{y}$ axes.
Reward: Agents are rewarded using a reward function that motivates agents to form a connected network bridging the targets. Such reward signal combines three components:
Base connectivity reward: The base reward, $R_{\text{base}}$, encourages larger connected components:
$R_{\text{base}}(s)=\frac{|C_{\text{max}}(s)|}{|\mathcal{V}|},$ (17)
where $|C_{\text{max}}(s)|$ represents the size of the largest connected component in state $s$, and $|\mathcal{V}|$ is the total number of entities.
Centroid distance penalty: The centroid distance penalty, $P_{\text{cent}}$, is based on agents’ positions relative to the targets, penalizing them based on the Euclidean distance between the centroid of the agents’ positions and the central point between the targets.
Target path bonus: A bonus, $B_{\text{path}}=100$, is awarded if a path exists between the two targets.
The overall reward combines these three elements:
$R(s,a)=\begin{cases}B_{\text{path}}(s)&\mbox{if }\exists\mathrm{path}(T_{1},T_{2});\par\\ R_{\text{base}}(s)-P_{\text{cent}}(s),&\mbox{otherwise}.\end{cases}$ (18)
Simulated UAV dynamics: For our experiments, we use a six-degrees-of-freedom (6-DOF) unmanned aerial vehicle (UAV) as an agent, where its dynamics, control laws, and safety sets used in our simulations are defined in the Appendix of [7]. For the UAV simulations, the nonlinear model of the drone is used.
Target mobility: To increase task complexity, each target moves deterministically toward waypoints located outside the communication range of any agent. When a target reaches a waypoint, it pauses for a fixed number of time-steps (5 in our implementation) before selecting a new one. Waypoint selection ensures that at least two agents (and up to $N$) are required to maintain a communication bridge between the targets.
Figure 3: A typical configuration of our system. Agents are represented with letters A, B, C and targets with T. Agent A and B have established a successful connection. Agent A receives node features $f_{n}^{A}$, $f_{n}^{A}$ and edge features $f_{n}^{AB}$ as inputs.
V-B Training configurations
We train our agents using episodes of 100 time steps, for a total of one million steps. For generalization, we initialize agent positions, target locations, and target motion patterns using different seeded random states across training and testing scenarios. During evaluation, each trained agent is tested in 100 new episodes, and we measure performance metrics with the safety filter consistently enabled.
We systematically evaluated different training configurations to assess the impact of our safety mechanism:
  • Baseline: A vanilla MARL approach with no safety filter enabled. Edge features include only the base features $\mathbf{f}$.
  • Approach A (Safety Filter): Our main contribution. Agents use Algorithm 1 to prevent safety infringements. The binary safety activation indicator is incorporated into the agents’ edge-level features ($\mathbf{f}$ and $S_{i,j}$).
  • Approach B (Safety + Penalty): We add an explicit penalty of $-10$ to Approach A whenever an agent attempts an action that would trigger the safety filter.
  • Approach C (Safety + Penalty + Early Truncation): Building on Approach B, we truncate episodes immediately upon safety-triggering situations, thus enforcing even stricter safety adherence.
Every approach adopts the same learning algorithm and GNN architecture. In any configuration employing the safety filter, an agent violating (7) has its trajectory toward $x_{sp}$ blocked at its current position.
Furthermore, we train an alternative of Approach A, called Approach A-Node, to investigate the benefits of our proposed integration of safety-informed edges IV-B. In this alternative approach, the agents utilize only the node-level features with the addition of a binary indicator $\mathbb{S}$ in case the agent had his safety filter activated during the previous step.
VI Results
(a) Mean return achieved during training. Approach A (with safety-informed edges) converges to higher rewards compared to Approach A-Node.
(b) Average number of safety filter activations per episode. Approach A consistently triggers fewer interventions.
Figure 4: Impact of safety-informed edge features.
Table I summarizes our experimental evaluation of various learned MARL strategies under different training conditions. This comparison was conducted to evaluate the impact of our safety filter. We report the average communication coverage achieved by the agents and the average number of times the safety filter was activated during evaluation.
Furthermore, Figure 4 focuses on a comparison between Approach A and Approach A-Node in terms of the mean return, i.e. the sum of the rewards, achieved by the agents during training and the mean number of safety activations triggered by the agents.
VI-A Impact of the safety filter
Baseline agents, trained without any safety mechanisms, achieved an average coverage of $(37.63\pm 16.62)\%$ but required frequent interventions from the safety filter, averaging $26.99\pm 12.31$ triggers per episode. In contrast, our Approach A, which incorporates our safety filter during training without explicit penalties, significantly improved average coverage to $(50.14\pm 12.72)\%$ and reduced safety filter activations to $20.45\pm 7.49$.
Introducing explicit penalties alongside the safety mechanism (Approach B) maintained competitive coverage at $(45.02\pm 11.81)\%$ and further reduced the frequency of safety filter interventions to $15.25\pm 7.82$. However, implementing truncation upon safety violations (Approach C) considerably decreased task performance to $(10.86\pm 6.44)\%$, though it also minimized the safety triggers to $9.76\pm 5.72$.
These results demonstrate that integrating our safety filter during training and including safety-informed edges allows agents to learn better strategies for the dynamic network bridging environment.
Results on Approach B and Approach C highlight how the introduction of explicit penalties might reduce safety interventions, which could have practical benefits in specific applications. However, this led to a decrease in task performance, which was conspicuous for Approach C, suggesting that agents might exploit the activation of the safety filter to achieve better performance.
VI-B Impact of the safety-informed edges
Figure 5: Sequence of actions executed by agents trained with Approach A when deployed in a LVC environment.
Figure 4(a) shows the mean return (cumulative reward) curves for Approach A and Approach A-Node, over the course of training. Approach A (with safety edge features) converges to higher returns compared to Approach A-Node, indicating that safety-informed edges facilitate more effective coordination strategies. Additionally, Figure 4(b) presents the average frequency of safety filter activations across training episodes. The results show that Approach A triggers fewer safety interventions than Approach A-Node, particularly in the later stages of training.
These findings suggest that edge-level safety features, rather than node-level, help achieve better performance while incurring in less safety activations. Figure 5 shows an example of a sequence of actions operated by our decentralized agents trained with Approach A when deployed in a live, virtual, and constructive (LVC) environment.
VII Discussion
Our findings demonstrate that integrating a decentralized safety filter with MARL provides a versatile solution for multi-agent networked tasks, where agents can directly communicate with their one-hop neighbors. By embedding the safety mechanism in each agent, we ensure local checks on potential collisions, thereby containing safety concerns within these immediate communication ranges. This design preserves the decentralized nature of our approach while providing global safety assurances. We applied the framework to the dynamic network bridging task, which we extended to include more realistic target mobility and non-linear UAV dynamics.
To support coordination, we included safety-informed edge features in the agents’ observations. These features indicate when the safety filter is active, allowing agents to adjust their actions based on local constraints. This helped agents avoid collisions, maintain communication links, and resolve potential deadlock situations through coordination.
Empirical evaluations confirm that our approach effectively improves task performance while assuring safety. Although adding explicit penalties or early-termination strategies reduced safety activations, these methods often compromised the agent’s utility. In contrast, our decentralized filter and safety-informed edges guide safe exploration without requiring extensive reward engineering or outcome truncations. It is important to highlight that the design principles generalize naturally to larger numbers of agents and targets and it can be applied to different environments than the one under study here. For example, we can apply similar safety filters to other networked navigation tasks or expand the existing safety filter to include complex metrics.
Looking ahead, we plan to investigate scalability for swarms of varying sizes, refine message-passing protocols for more sophisticated risk management, and evaluate domains with adversarial behaviors or dynamic constraints. Overall, our results highlight the potential of combining local safety controllers with MARL in networked multi-agent settings, underscoring that edge-level safety signals are crucial for guiding agents toward globally safe and high-performing outcomes.